Kindly assist me solving the following questions
1. (x^2y^2+xy)ydx + (x^2y^2-1)xdy=0
2. (x^2+y^2)dx + xdy=0
3. dy/dx +y=3y^3e^x
1
Expert's answer
2019-08-13T11:30:19-0400

1.


"xy^2(xy+1)dx+x(xy-1)(xy+1)dy=0"

"x=0""y=0""xy+1=0=>y=-{1 \over x}"

Or


"y^2dx+(xy-1)dy=0"

Using an Integrating Factor

"P(x,y)=y^2, Q(x,y)=xy-1"

"{\eth P \over \eth y}=2y, {\eth Q \over \eth x}=y, {\eth P \over \eth y}\not={\eth Q \over \eth x}""{\eth P \over \eth y}-{\eth Q \over \eth x}=2y-y=y"

Integrating Factor

"\mu=\mu(y)"

"{1 \over \mu}{d\mu \over dy}=-{1 \over P}({\eth P \over \eth y}-{\eth Q \over \eth x})"

"{1 \over \mu}{d\mu \over dy}=-{1 \over y^2}(2y-y)"

"{d\mu \over \mu}=-{dy \over y}"

"\int {d\mu \over \mu}=\int (-{dy \over y})=>\ln|\mu|=\ln|{1 \over y}|+\ln C"

We choose


"\mu={1 \over y}"

Multiplying the original differential equation by "\mu=1/y"  produces the exact equation:


"ydx+(x-{1 \over y})dy=0"

Indeed, now we have


"{\eth P \over \eth y}=1, {\eth Q \over \eth x}=1, {\eth P \over \eth y}={\eth Q \over \eth x}"

Solve the resulting equation. The function "u(x,y)" can be found from the system of equations:


"\begin{cases}
 {\eth u \over \eth x}=y \\
 {\eth u \over \eth y}=x-{1 \over y}
\end{cases}"

It follows from the first equation that


"u=\int y dx=xy+\varphi (y)"

Substitute this in the second equation to determine "\varphi:"


"{\eth u \over \eth y}=x+ {d \varphi \over d y}=x-{1 \over y}"

"\int d\varphi=\int (-{1 \over y})dy=>\varphi (y)=\ln({1 \over |y|})"

Thus, the original differential equation has the following solutions:


"x=0""y=0""y=-{1 \over x}""xy+\ln({1 \over |y|})=C_1"

where "C_1" is a constant.


2.


"(x^2+y^2-y)dx+xdy=0"

We can make sure that this equation is not exact:


"P(x,y)=x^2+y^2-y, Q(x,y)=x"

"{\eth P \over \eth y}=2y-1, {\eth Q \over \eth x}=1, {\eth P \over \eth y}\not={\eth Q \over \eth x}"

The difference of the partial derivatives is


"{\eth P \over \eth y}-{\eth Q \over \eth x}=2y-1-1=2y-2"

Using the integrating factor "\mu=x^2+y^2" we find that


"{\eth z \over \eth y}=2y, {\eth z \over \eth x}=2x"

Calculate the following expression:

"Q{\eth z \over \eth x}-P {\eth z \over \eth y}=x(2x)-(x^2+y^2-y)(2y)=""=2x^2+2y^2-2y(x^2+y^2)=(x^2+y^2)(2-2y)"

As a result, we obtain the differential equation for the function "\mu(z):"

"{1 \over \mu}{d\mu \over dz}={1 \over (Q{\eth z \over \eth x}-P {\eth z \over \eth y})}({\eth P \over \eth y}-{\eth Q \over \eth x})="

"={1 \over (x^2+y^2)(2-2y)}(2y-2)=-{1 \over {x^2+y^2}}=-{1 \over z}"

By integrating we get the function "\mu(z):"

"\int {d\mu \over \mu}=\int (-{dz \over z})=>\ln|\mu|=-\ln|z|+\ln C"

We choose

"\mu={1 \over z}"

We can choose the integrating factor

"\mu=-{1 \over z}=-{1 \over {x^2+y^2}}"

"-{1 \over {x^2+y^2}}(x^2+y^2-y)dx-{1 \over {x^2+y^2}}xdy=0"

Or

"(1-{y\over {x^2+y^2}})dx+{x \over {x^2+y^2}}dy=0"

The general solution "u(x,y)=C" is defined by the following system of equations:

"\begin{cases}
 {\eth u \over \eth x}=1-{y\over {x^2+y^2}} \\
 {\eth u \over \eth y}={x \over {x^2+y^2}}
\end{cases}"

Integrating the first equation with respect to x produces:

"u(x,y)=\int{(1-{y\over {x^2+y^2}})}dx=x-y\int{{1\over {x^2+y^2}}}dx="

"=x-\arctan({x \over y})+\varphi(y)"

Substituting this in the second equation, we have:


"{\eth u \over \eth y}={x\over {x^2+y^2}}+ {d \varphi \over d y}={x\over {x^2+y^2}}"

"{d \varphi \over d y}=0"

Thus, the general solution of differential equation in implicit form is defined by the formula:

"x-\arctan({x \over y})=C"

where "C" is a constant.



3.


"{dy \over dx}+y=3y^3e^x""y=0"

Or


"y^{-3}{dy \over dx}+y^{-2}=3e^x"

The associated homogeneous differential equation is


"y^{-3}{dy \over dx}+y^{-2}=0"

"{dy \over dx}+y=0"

"{dy \over y}=-dx""y=C_1e^{-x}"

"{dy \over dx}={dC_1 \over dx}e^{-x}-C_1e^{-x}"

Substitute


"{dC_1 \over dx}e^{-x}-C_1e^{-x}+C_1e^{-x}=3C_1^3e^{-3x}e^x"

"{dC_1 \over dx}=3C_1^3e^{-x}"

"{dC_1 \over C_1^3}=3e^{-x}dx"

"\int {dC_1 \over C_1^3}=\int3e^{-x}dx"

"-{1 \over2 C_1^2}=-3e^{-x}dx-{C_2 \over2}"

"C_1={1\over \sqrt{6e^{-x}dx+C_2}}"

Thus, the original differential equation has the following solutions:


"y=0""y={1\over e^x \sqrt{6e^{-x}dx+C_2}}"

where "C_1" is a constant.